you normally drive on the freeway between butuan city and cagayan de oro city at an average speed of 100km/hr (62.14 mi/hr), and the trip takes 4 hours and 30 mins on a friday afternoon, however, heavy traffic slows you down and you drive the same distance at an average speed of only 70km/hr (43 mi/hr). how much longer does the trip take?


1
Expert's answer
2022-02-19T04:55:49-0500

For the case of a normal trip, we determine the distance between cities

We have

"v=100km\/hr"

"t= 4hr30mins=4.5 hr"

then the distance between cities is

"S=v \cdot t=100 \cdot 4.5 = 450km"

then for the average speed

"u=70 km\/hr"

travel time is

"\tau=\frac{S}{u}=\frac{450}{70}=6.43hr=6hr25mins48sec"